OpenStudy (anonymous):

Emily plans to use the quadratic formula, while Juan is considering using the graphing method to find exact solutions of 3x^2 – 8x = 20. Using complete sentences, explain who has chosen the better method and why. Emily plans to use the quadratic formula, while Juan is considering using the graphing method to find exact solutions of 3x^2 – 8x = 20. Using complete sentences, explain who has chosen the better method and why. @Mathematics

hero (hero):

actually, I take that back. Use the quad formula for this. If you tried to graph it, it wouldn't give you any "solutions"

hero (hero):

At least with the quadratic formula, you can still get the imaginary solutions

OpenStudy (anonymous):

Thank you.

OpenStudy (radar):

I don't see any imaginary solutions, maybe some irrational solutions.\[3x ^{2}-8x-20=0\]\[8\pm \sqrt{(-8)^{2}-4(3)(-20)}\over6\]\[8\pm \sqrt{64+240}\over6\]\[8\pm \sqrt{304}\over6\]\[8\pm4\sqrt{19}\over6\]\[4\pm2\sqrt{19}\over3\] Looks like one positive irrational root and one negative irrational root. No imaginary roots as the value within the radical is not negative.

OpenStudy (radar):

However, quadratic is the way to go, because it is hard to plot those irrational values lol.